The following production data were taken from the records of the Finishing Department for June:

Inventory in process, June 1 (30% completed) 4,000 units

Completed and transferred units during June                                                   65,000 units

Ending inventory (60% complete) 10,000 units

What is the cost per conversion equivalent units of production at June 30, assuming that the weighted average method is used to cost inventories and total conversion costs were $122,000?

1.20

1.72

1.63

1.54
Answer #1

Equivalent unit of production = 65000+(10000*60%) = 71000 Units

Cost per equivalent unit = Total conversion Cost/Equivalent unit = 122000/71000 = 1.72

So answer is b) $1.72